# Generalizable Deepfake Detection via Effective Local-Global Feature Extraction

### v1

@@ 3.1 Window Attention Block
3.2 DWT-Based Local Spatial-Frequency Domain Feature Capture Block
Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain. However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features. To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$. As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $2\times 2$ windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the $2\times 2$ windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $2\times 2$ window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$. To further enhance the network’s performance, we apply LayerNorm Lei Ba et al. (2016), which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
The feature ${\hat{I}}_{\text{input}}$ is then fed into WAttention with a sliding window size set to $b=4$. As depicted in Figure 3, within each $4\times 4$ window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands. This configuration provides WAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction. Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
3.3 FFT-Based Global Frequency Domain Feature Capture Block
Although artifact information is often concentrated in finer details, forged features can also be embedded in the global information of an image, remaining imperceptible to the naked eye. While DWT and sliding window tiling are effective for extracting local features, we also use FFT capture global frequency domain information from a complementary perspective. In this module, FFT is employed to transform the image into the frequency domain at the pixel level, enabling the extraction of global frequency domain features. Inspired by Zhao et al. (2024), we recognize that amplitude primarily reflects properties such as color and brightness, while phase encodes finer details, including potential artifacts. Based on this insight, we focus on extracting the phase features as input to WAttention, setting the window size to $b=8$. The amplitude remains unmodified during this process. After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain. The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
Finally, we set a hyperparameter $\lambda$ to minimize the impact of global information extraction. The final feature is:
Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$. In the subsequent LoAttention module, $4\times 4$ features are extracted as shown in the blue box, which not only contains local spatial frequency information, but also contains all different frequency band features.
4 Experiments
To verify the excellent deepfake detection performance of our method, we followe the paradigms of Ojha et al. (2023) and Wang et al. (2020).
4.1 Dataset
Training set. To ensure a consistent basis for comparison, we use the training set of ForenSynthsWang et al. (2020). The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset. Following the paradigms of Ojha et al. (2023), we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
Testing set. In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion datasets for testing, which contain a wide range of real-world images.
  • ForenSynths: The test set comprises 10 subsets, including: ProGAN Karras (2017), StyleGAN Karras (2019), StyleGAN2 Karras et al. (2020), BigGAN Brock (2018), CycleGAN Zhu et al. (2017), StarGAN Choi et al. (2018), GauGAN Park et al. (2019), Deepfake Rossler et al. (2019), WFIR and SAN Dai et al. (2019).
  • GANGen-Detection Tan et al. (2024a): To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images. There are 4K test images for each model, with equal numbers of real and fake images.
  • DiffusionForensics Wang et al. (2023): To expand the testing scope, we adopt the diffusions dataset of DIRE Wang et al. (2023) for evaluation, including ADM Dhariwal and Nichol (2021), DDPM Ho et al. (2020), IDDPM Nichol and Dhariwal (2021), LDM Rombach et al. (2022), PNDM Liu et al. (2022), VQ-Diffusion Gu et al. (2022), Stable Diffusion v1 Rombach et al. (2022), Stable Diffusion v2 Rombach et al. (2022). The real images are sampled from LSUN Yu et al. (2015) and ImageNet Russakovsky et al. (2015) datasets.
  • UniversalFakeDetect Ojha et al. (2023): This test set contains images generated from Glide Nichol et al. (2021), DALL-Emini Ramesh et al. (2021), LDM Rombach et al. (2022). It adopts images of LAION Schuhmann et al. (2021) and ImageNetRussakovsky et al. (2015) datasets as the real data.
  • AIGCDetectBenchmark Zhu et al. (2024): This dataset mainly adopts diffusion model to generate images, and is generated using more realistic advanced methods, including Midjourney, SDv1.4, SDv1.5 Rombach et al. (2022), ADM Dhariwal and Nichol (2021), GLIDE Nichol et al. (2021), Wukong, VQDM Gu et al. (2022), which is a huge challenge for deepfake detection.
4.2 Implementation Details
We design a lightweight CNN network using convolutional layers and Resnet network as classifiers. The network is trained using the Adam optimizerKingma (2014) with a learning rate of $2\times 10^{-4}$. The batchsize is set to 128, and we train the network for 90 epochs. We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs. What’s more, the hyperparameter $\lambda$ is set to 0.2. Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU. To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines Ojha et al. (2023); Tan et al. (2024a), which include average precision score (A.P.) and accuracy (Acc.).
4.3 Quantitative analysis
We compare with the previous methods: CNNDetection Wang et al. (2020), Patchfor Chai et al. (2020), F3Net Qian et al. (2020), FrePGAN Jeong et al. (2022), GANDetection Mandelli et al. (2022), LGrad Tan et al. (2023), Ojha Ojha et al. (2023), FreqNet Tan et al. (2024a) and NPR Tan et al. (2024b). We conduct two and three sets of experiments on GAN and diffusion respectively.
GAN. As shown in Table 1 and 2, our detection on GAN has achieved good results. Table 1 presents a detailed comparison of detector performance on the test set of ForenSynths Wang et al. (2020), highlighting higher mean accuracy (Acc.) and comparable mean average precision (A.P.) metrics. For unavailable test results, we use ”-” as a placeholder and exclude these values from the average calculations. Our method achieves an average accuracy of 86.9%, which is 1% and 1.7% higher than the latest methods, NPR and FreqNet, underscoring its effectiveness in generalizable deepfake detection. Furthermore, to evaluate cross-GAN-source generalization, we conducted tests on the GANGen-Detection dataset Tan et al. (2024a), as summarized in Table 2. Our method achieves an impressive average accuracy of 96.5%, significantly outperforming the best-performing baselines, NPR Tan et al. (2024b) and FreqNet Tan et al. (2024a), which achieved 94.8% and 94.0%, respectively.
Diffusion. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables 3 and 4. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.5% and 99.9%, respectively. Compared to NPR and FreqNet, our method improves mean Acc. by 2.9% and 17.8%, respectively. Additionally, on the UniversalFakeDetect dataset Ojha et al. (2023), our method achieves an average accuracy of 95.0%, representing a notable improvement over existing approaches.
With the rapid advancements in tools like Midjourney and diffusion models, increasingly realistic fake images are being generated, posing significant challenges for deepfake detection. To address this, we conduct additional evaluations on the AIGCDetectBenchmark dataset Zhu et al. (2024), excluding some subsets of data previously tested. The results, summarized in Table 5, highlight the robustness of our approach. Our method achieves a notable improvement in mean accuracy, outperforming NPR and FreqNet by 3.7% and 7.8%, respectively, and reaching an accuracy of 77.6%.
Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR SAN Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 91.4 99.4 63.9 92.2 76.4 97.5 52.9 73.3 63.8 91.4 72.7 88.6 63.8 92.2 51.7 62.3 91.7 96.3 48.7 59.5 67.7 85.3
Patchfor 97.8 100.0 57.2 55.4 83.6 98.5 64.7 69.5 82.6 93.1 74.5 87.2 100.0 100.0 85.0 93.2 - - 75.3 76.3 80.1 85.9
F3Net 99.4 100.0 58.1 56.1 88.0 99.8 65.3 69.9 92.6 99.7 76.4 84.3 100.0 100.0 63.5 78.8 - - 47.3 46.7 76.7 81.7
FrePGAN 99.0 99.9 60.3 77.1 84.1 98.6 69.2 71.7 80.7 89.6 71.1 74.4 99.9 100.0 70.9 91.9 - - - - 79.3 87.9
LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1 50.0 51.4 79.9 84.8
Ojha 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7 56.5 82.2 84.2 92.6
FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6 71.9 80.1 85.2 91.4
NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.6 62.9 64.4 66.0 85.9 87.1
ours 99.9 100.0 86.3 97.4 99.9 100.0 87.1 93.6 99.9 100.0 97.1 98.9 90.8 95.9 64.5 72.2 77.8 87.9 65.6 70.6 86.9 91.7
Table 1: Cross-GAN-Sources Evaluation on the test set of ForenSynths Wang et al. (2020). Bold and underline represent the best and second-best performance, respectively.
@@ 4.3 Quantitative analysis
ours 97.7 99.8 96.4 99.6 98.2 99.8 98.8 99.9 98.8 88.8 99.4 100.0 83.4 90.2 98.7 99.8 96.8 99.7 96.5 98.7
Table 2: Cross-GAN-Sources Evaluation on the 9 GANs dataset Tan et al. (2024a).
@@
NPR 99.4 100.0 100.0 100.0 97.3 99.8 88.7 98.6 96.4 100.0 87.5 98.3 97.5 100.0 100.0 100.0 95.8 99.6
ours 99.4 100.0 99.9 100.0 95.6 99.7 96.3 99.4 99.2 100.0 99.4 100.0 99.4 100.0 99.8 100.0 98.7 99.9
Table 3: Cross-Diffusion-Sources Evaluation on the test of DiffusionForensics Wang et al. (2023).
@@
NPR 74.8 78.8 98.3 99.8 98.7 99.7 98.2 99.8 98.8 99.8 99.0 99.8 98.8 99.7 95.9 99.7 94.8 97.0
ours 82.4 87.1 98.3 99.7 97.6 99.8 97.6 99.6 99.4 100.0 99.3 100.0 99.0 99.9 86.0 95.6 95.0 97.7
Table 4: Cross-Diffusion-Sources Evaluation on the diffusion test set of Ojha Ojha et al. (2023).

### v2

@@ III-B Reconstructed Sliding Window Attention Block
TABLE I: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the GANGen-Detection Dataset [9].
@@
FatFormer 99.9 100.0 98.3 99.9 98.4 100.0 98.4 100.0 99.5 100.0 98.8 99.8 99.0 100.0 99.3 100.0 98.4 100.0 98.9 100.0
Ladeda 100.0 100.0 99.0 99.9 97.9 99.9 99.0 99.9 100.0 100.0 99.8 100.0 80.5 84.9 100.0 100.0 98.4 99.9 97.2 98.3
C2P_CLIP 94.9 100.0 98.4 99.9 98.4 100.0 98.4 100.0 92.0 99.8 97.6 99.6 99.0 100.0 90.4 99.8 98.4 100.0 96.4 99.9
AIDE 57.9 79.5 76.1 79.2 82.0 88.1 82.2 88.3 95.2 96.5 92.7 98.1 69.0 78.8 96.6 98.4 80.7 82.8 81.4 87.8
Ours 98.6 99.7 96.8 99.6 98.0 99.7 98.7 99.8 99.0 99.9 99.1 100.0 88.2 91.1 99.0 99.8 97.2 99.7 97.3 98.8
TABLE II: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the DiffusionForensics Dataset[46].
@@
NPR 99.4 100.0 100.0 100.0 97.1 99.7 88.4 97.9 96.4 100.0 87.5 98.0 97.5 100.0 100.0 100.0 95.8 99.4
FatFormer 67.2 73.5 100.0 100.0 61.7 96.8 70.8 93.4 84.4 98.2 69.3 94.3 99.3 100.0 97.3 100.0 81.2 94.4
Ladeda 98.4 99.9 100.0 100.0 97.1 99.8 83.9 96.9 94.1 100.0 93.9 98.8 93.7 99.9 100.0 100.0 95.1 99.4
C2P_CLIP 73.4 76.2 95.8 99.7 78.9 99.2 68.8 95.3 66.7 94.8 80.7 94.9 84.2 97.2 97.3 99.7 80.7 94.6
AIDE 86.5 92.5 92.1 99.4 91.5 98.6 85.7 94.1 88.6 95.1 87.8 94.8 89.6 97.8 87.4 95.0 88.6 95.9
Ours 99.9 100.0 100.0 100.0 97.5 99.9 96.0 99.8 99.9 100.0 98.5 100.0 99.2 100.0 99.6 100.0 98.8 100.0
III-C Dual Frequency Branch Feature Enrichment Module
DWT-based window tiling branch. Previous methods often leverage the Fast Fourier Transform (FFT) or Discrete Cosine Transform (DCT) to capture relationships between adjacent nodes in the frequency domain. However, these approaches can lead to the loss of positional information, which is crucial for extracting detailed features. To overcome this limitation, we utilize the Discrete Wavelet Transform (DWT), which decomposes the image into the spatial-frequency domain, represented as:
$\text{DWT}(I)=\{I_{LL},I_{LH},I_{HL},I_{HH}\},$ (4)
where $I_{LL},I_{LH},I_{HL},I_{HH}\in{\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}}$ represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features $\hat{I}\in{\mathbb{R}^{4\times C\times\frac{H}{2}\times\frac{W}{2}}}$. As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into $\times 2$ 2 windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the $\times 2$ 2 windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the $\times 2$ 2 window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features ${\hat{I}}_{\text{input}}\in{\mathbb{R}}^{C\times 4\times\frac{H\times W}{4}}$. To further enhance the network’s performance, we apply LayerNorm [47], which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
The feature ${\hat{I}}_{\text{input}}$ is then fed into RSWAttention with a sliding window size set to $b=4$. As depicted in Figure 3, within each $\times 4$ 4 window, the representation includes not only the four adjacent feature values from the original image but also the feature information across four distinct frequency bands. This configuration provides RSWAttention with comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction. Following this process, we apply the IDWT to reconstruct the features back into spatial domain representations, denoted as ${\hat{I}_{\text{DWT\_output}}}$.
FFT-based phase complement branch. Different from DWT, which captures local spatial frequency features, each element of the phase part in the FFT domain contains the dependencies and structural relationship between the overall elements, which helps the model capture the long-range dependencies between features. In addition, through our experimental analysis, we found that phase encodes finer details, including potential artifacts. Therefore, in order to enrich the extraction of local fine-grained features from another perspective, we only use the phase part of the FFT as another branch input of the RSWAttention, setting the window size to $b=8$. The amplitude remains unmodified during this process. After training the phase features, we use IFFT to recombine them with the original amplitude, transforming the features back into the spatial domain. The resulting features are denoted as ${\hat{I}_{\text{FFT\_output}}}$.
Finally, we set a hyperparameter $\lambda$ to effectively combines the outputs of two branches. The final feature is:
TABLE III: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the UniversalFakeDetect Dataset [43].
@@ III-C Dual Frequency Branch Feature Enrichment Module
NPR 74.0 78.1 97.5 99.5 97.8 99.5 97.4 99.5 98.2 99.6 98.2 99.6 98.0 99.5 90.9 98.1 93.9 96.7
FatFormer 76.1 92.0 94.7 99.4 94.2 99.2 94.4 99.1 98.7 99.9 98.6 99.8 94.9 99.1 98.8 99.8 93.8 98.5
Ladeda 79.8 87.5 98.3 99.8 98.2 99.8 98.5 99.8 98.6 99.9 98.5 99.9 98.0 99.8 83.9 98.5 94.2 98.1
C2P_CLIP 69.1 94.1 95.3 99.8 96.1 99.8 95.3 99.8 99.3 100.0 99.3 100.0 97.3 99.8 98.6 99.1 93.8 99.2
AIDE 80.4 92.6 94.6 98.9 94.7 98.8 94.5 98.8 93.1 98.6 93.0 98.3 93.3 98.6 92.2 98.0 91.9 97.8
Ours 95.2 97.8 98.0 99.9 97.7 99.9 97.2 99.7 99.4 100.0 99.5 100.0 98.9 99.9 85.2 97.1 96.4 99.3
TABLE IV: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the GenImage Datasets [48].
Method ProGAN GauGAN StyleGAN2 BigGAN StyleGAN CycleGAN StarGAN Deepfake WFIR
CNNDetection 97.9 99.9 66.1 93.5 67.0 97.1 55.9 80.5 70.1 97.0 76.5 91.5 72.9 93.2 51.7 66.5 62.1 89.4
LGrad 99.9 100.0 72.4 79.3 96.0 99.9 82.9 90.7 94.8 99.9 85.3 94.0 99.6 100.0 58.0 67.9 60.5 65.1
CLIPDetection 100.0 100.0 99.7 100.0 75.7 97.8 95.1 99.3 84.4 97.1 98.7 99.8 99.9 99.4 67.4 82.0 64.2 68.7
FreqNet 99.6 100.0 93.4 98.6 88.0 99.5 90.5 96.0 90.2 99.7 95.8 99.6 85.7 99.8 88.9 94.4 48.0 49.6
NPR 99.9 100.0 80.9 83.0 99.6 100.0 84.0 85.6 98.1 99.8 95.2 98.1 99.8 100.0 77.2 76.0 60.7 63.9
FatFormer 99.9 100.0 99.4 100.0 98.8 99.9 99.5 99.9 97.2 99.8 99.3 100.0 99.8 100.0 93.2 98.0 88.2 98.5
Ladeda 100.0 100.0 94.1 100.0 99.9 100.0 82.5 89.5 99.8 100.0 91.1 99.0 76.2 90.3 64.0 92.6 85.7 93.6
C2P_CLIP 100.0 100.0 99.2 100.0 95.6 99.9 99.1 100.0 96.4 99.5 97.3 100.0 99.6 100.0 93.8 98.6 94.8 99.5
AIDE 98.0 99.8 60.2 71.9 97.5 99.6 71.0 82.2 97.4 99.5 86.0 97.1 99.3 100.0 53.7 70.3 90.3 97.0
Ours 99.8 100.0 85.1 96.7 99.9 100.0 87.1 95.1 99.8 100.0 97.9 98.8 94.9 99.2 64.2 71.8 52.2 46.8